150 question - each 1 mark for wright answer
if wrong - 1 mark will cut for every 3 question
i got 90 marks then how much correct answer did i got

Answer by srinivas.g(540)

let x be the no of correct answers you did
let y be the no of wrong answers you did
so x+y = 150 ----eq(1)
marks you got is 90
so x*(1)+y(-1)= 90
x-y =90
move y to the right
x = 90+y
put this x= 90+yin x+y= 150
(90+y)+y= 150
90+2y = 150
subtract 90 on both sides
90+2y -90 =150-90
2y = 60
divide with 2 on both sides
2y/2 = 60/2
y= 30
x= 90+y= 90+30 =120
so no of correct answers you made = 120